OpenStudy (anonymous):

write algebraic expression w/0 partenthesis -4(-5x)=

OpenStudy (anonymous):

would this answer be -4 *-5x=

OpenStudy (lgbasallote):

yes. but you have to simplify it furhter

OpenStudy (lgbasallote):

further*

OpenStudy (anonymous):

Yes multiply that ..

OpenStudy (lgbasallote):

what is the value of \[-4 \times -5\]??

OpenStudy (anonymous):

can you multiple a number without a (x) with a x number

OpenStudy (lgbasallote):

yes you can. but you cannot add them

OpenStudy (lgbasallote):

\[a \times bx \implies (a \times b)x\]

OpenStudy (anonymous):

For example: -3*-4x = (-3*-4)x = 12x

OpenStudy (anonymous):

for some reason i got -185 i was putting ((-4x)-5) in my caculator

OpenStudy (lgbasallote):

just do \[-4 \times -5\] then just stick the x later on

OpenStudy (anonymous):

There is no need of calculator here..

OpenStudy (lgbasallote):

how are you getting -185? o.O

OpenStudy (anonymous):

What did you get when you multiply this : -4*(-5) = ??

OpenStudy (anonymous):

i have took this clas 5 times i am beyond bad in math this class is the only thing keeping me from finishing college. Dont hate on me using a caculator i cant even factor

OpenStudy (lgbasallote):

it's alright if you use a calculator. we all do.

OpenStudy (lgbasallote):

calculators were invented to be used after all

OpenStudy (anonymous):

when i put in -4(-5) i got 20

OpenStudy (lgbasallote):

that's right

OpenStudy (lgbasallote):

now attach an x

OpenStudy (anonymous):

so its 20x

OpenStudy (lgbasallote):

yes

OpenStudy (anonymous):

Yep..

OpenStudy (anonymous):

ok thank you for your help